OpenStudy (anonymous):

I'm trying to use integrals to find the arc length of y= ln(secx), 0

OpenStudy (anonymous):

I understand the formula but Im kind of confused on how to use it on this particular problem

OpenStudy (anonymous):

Arc Length = ∫ (√(1 + [f '(x)]^2)) dx {from a to b} ∫ (√(1 + [tan(x)]^2)) dx {from 0 to pi/4} Use the trig identity sec(x)^2 = 1 + tan(x)^2 = ∫ √(sec(x)^2) dx {from 0 to pi/4} = ∫ sec(x) dx {from 0 to pi/4} = ∫ 1/cos(x) dx {from 0 to pi/4} Multiply the integrand by cos(x)/cos(x). = ∫ cos(x) / cos(x)^2 dx {from 0 to pi/4} = ∫ cos(x) / (1 - sin(x)^2) dx {from 0 to pi/4} Now do a substitution. Let u=sin(x), then du=cos(x) dx. ∫ cos(x) / (1 - sin(x)^2) dx {from 0 to pi/4} = ∫ du / (1 - u^2) {from 0 to pi/4} = ∫ du / ((1 - u)*(1 + u)) {from 0 to pi/4} = (1/2) * ∫ (1 / (1 - u) + 1 / (1 + u)) du {from 0 to pi/4} = (1/2) * [∫ (1 / (1 - u) du + ∫ 1 / (1 + u)) du] {from 0 to pi/4} = (1/2) * [ - ln(|1 - u|) + ln(|1 + u|) ] {from 0 to pi/4} = ln(|1 + u|) / 2 - ln(|1 - u|) / 2 {from 0 to pi/4} = ln(|1 + sin(x)|) / 2 - ln(|1 - sin(x)|) / 2 {from 0 to pi/4} = (ln(|1 + sin(pi/4)|) / 2 - ln(|1 - sin(pi/4)|) / 2) - (ln(|1 + sin(0)|) / 2 - ln(|1 - sin(0)|) / 2) = ln(|1 + sin(pi/4)|) / 2 - ln(|1 - sin(pi/4)|) / 2 = ln(|1 + √2 / 2|) / 2 - ln(|1 - √2 / 2|) / 2 = [ln(|1 + √2 / 2|) - ln(|1 - √2 / 2|)] / 2 = ln(|(1 + √2 / 2) / (1 - √2 / 2)|) / 2 = ln(2*√2 + 3) / 2 = ln((2*√2 + 3)^(1/2)) The √(2*√2 + 3) = 1 + √2, so ln((2*√2 + 3)^(1/2)) = ln(1 + √2)

OpenStudy (anonymous):

but how did you get to the first step without getting 1/secx

OpenStudy (anonymous):

i have just go it. i cant explain u

OpenStudy (anonymous):

well ln(sec(x)) the derivative of this is 1/secx right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then you square it. and you get 1/sec^2x

OpenStudy (anonymous):

how did you get 1+tan^2x in the first step of you integral

OpenStudy (anonymous):

cos + x square

OpenStudy (anonymous):

where did you get cos + x square?